# On the splitting of weak nearly $\cal C$}-manifolds

### v1

@@ 3 Auxiliary results
This proves the first part, while the latter follows directly as a consequence of the commutativity and skew-symmetry of the operators $h_{i}\ (1\leq i\leq s)$. ∎
Lemma 7.
For a weak nearly $\cal C$}-manifold satisfying (5), (6), (9) and (13), we have
$g((\nabla_{X}{f})Y,{f}h_{j}Z)=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(X)g(h_{j}Y,h_{i}Z)-\eta^{i}(Y)g(h_{j}X,h_{i}Z)\big]\quad(1\leq j\leq s).$ (49)
Proof.
Differentiating (19) and using $g((\nabla_{X}{f})(\nabla_{V}{f})Y,Z)$ $g((\nabla_{V}{f})Y,(\nabla_{X}{f})Z)$ gives
$g((\nabla_{V}\,{f})Y,(\nabla_{X}{f})Z)+g((\nabla_{X}{f})Y,(\nabla_{V}\,{f})Z)=g((\nabla^{2}_{V,X}\,{f}){f}Y,Z)$
$+g((\nabla^{2}_{V,X}\,{f}){f}Z,Y)-\sum\nolimits_{i=1}^{s}\big[g(h_{i}V,Z)\,g(h_{i}X,Y)+\eta^{i}(Z)\,g((\nabla_{V}h_{i})X,Y)$
$+g(h_{i}V,Y)\,g(h_{i}X,Z)+\eta^{i}(Y)\,g((\nabla_{V}h_{i})X,Z)+\nabla_{V}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)\big].$ (50)
Using (1), (25) and (31), we find $\nabla^{2}$-terms in (3):
$g((\nabla^{2}_{V,X}\,{f}){f}Z,Y)=g((\nabla^{2}_{V,{f}Z}\,{f})Y,X)=-g(R_{X,Y}V,QZ)+\sum\nolimits_{i=1}^{s}\eta^{i}(Z)\,g(R_{X,Y}V,\xi_{i})$
$\qquad-(1/2)\,g(R_{X,V}{f}Z,{f}Y)-(1/2)\,g(R_{V,Y}{f}Z,{f}X)-(1/2)\,g(R_{X,Y}{f}Z,{f}V),$ (51)
$g((\nabla^{2}_{V,X}\,{f}){f}Y,Z)=g((\nabla^{2}_{V,{f}Y}\,{f})Z,X)=-g(R_{X,Z}V,QY)+\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,g(R_{X,Z}V,\xi_{i})$
$\qquad-(1/2)\,g(R_{X,V}{f}Y,{f}Z)-(1/2)\,g(R_{V,Z}{f}Y,{f}X)-(1/2)\,g(R_{X,Z}{f}Y,{f}V).$ (52)
Substituting (3)–(3) in (3), we get
$-(1/2)\,g(R_{V,Z}{f}Y,{f}X)-(1/2)\,g(R_{X,Z}{f}Y,{f}V)-(1/2)\,g(R_{V,Y}{f}Z,{f}X)-(1/2)\,g(R_{X,Y}{f}Z,{f}V)$
$+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)\,g(R_{X,Z}V,\xi_{i})+\eta^{i}(Z)\,g(R_{X,Y}V,\xi_{i})-g(h_{i}V,Z)\,g(h_{i}X,Y)-\eta^{i}(Z)\,g((\nabla_{V}h_{i})X,Y)$
$-g(h_{i}V,Y)\,g(h_{i}X,Z)-\eta^{i}(Y)\,g((\nabla_{V}h_{i})X,Z)-\nabla_{V}\big(\eta(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)\big].$ (53)
$-g(R_{V,Z}{f}X,{f}Y)-g(R_{X,Z}{f}V,{f}Y)+\sum\nolimits_{i=1}^{s}\big[g(h_{i}V,Y)\,g(h_{i}X,Z)+g(h_{i}V,Z)\,g(h_{i}X,Y)\big]$
$=-(1/4)\delta(X,Z,V,Y)-(1/4)\delta(V,Z,X,Y)-\sum\nolimits_{i=1}^{s}\nabla_{V}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big).$ (54)
Replacing in (3) $Z$ by fZ{f}Z and $V$ by fV{f}V, we find
$+g(R_{X,Y}{f}V,{f}QZ)-g(R_{{f}V,{f}Z}{f}X,{f}Y)+\sum\nolimits_{i=1}^{s}\big[g(h_{i}X,Y)\,g(h_{i}{f}V,{f}Z)$
$+g(h_{i}X,{f}Z)\,g(h_{i}{f}V,Y)\big]=(1/4)\,\delta(X,{f}Z,{f}V,Y)-(1/4)\,\delta({f}V,{f}Z,X,Y).$ (55)
Using (1), (19), (20), (3) and Lemma 2, we find
$g((\nabla_{X}{f}){f}Z,(\nabla_{{f}V}\,{f})Y)=g(Q(\nabla_{X}{f})Z,(\nabla_{V}\,{f})Y)-\sum\nolimits_{i,j=1}^{s}\eta^{i}(Z)\,\eta^{j}(V)g(h_{i}X,h_{j}Y)$
$\quad+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Z)g({f}h_{i}X,(\nabla_{V}\,{f})Y)-g(X,{f}h_{i}Z)\,g(V,{f}h_{i}Y)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,{f}h_{i}Y)$
$\quad+g(h_{i}X,QZ)g(h_{i}V,QY)\big],$ (56)
$g((\nabla_{X}{f})Y,(\nabla_{{f}V}{f}){f}Z)=-g(Q(\nabla_{X}{f})Y,(\nabla_{V}\,{f})Z)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)g({f}h_{i}Z,(\nabla_{X}{f})Y)$
$\quad-\eta^{i}(Z)g({f}h_{i}V,(\nabla_{X}{f})Y)+g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)\big].$ (57)
From (45) and Lemma 2, we have
$g(R_{X,{f}Z}{f}V,Y)-g(R_{X,{f}Z}{f}^{2}V,{f}Y)$ $=g(R_{X,{f}Z}{f}V,Y)+g(R_{X,{f}Z}QV,{f}Y)$
$-\sum\nolimits_{i,j=1}^{s}\eta^{i}(X)\,\eta^{j}(V)\,g(h_{j}h_{i}Z,QY).$ (58)
On the other hand, from (22), (46) and (1) it follows that
$-g(R_{{f}X,Z}{f}V,QY)$ $+\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,g(R_{{f}X,Z}{f}V,\xi_{i})=-g(R_{QX,{f}Z}V,{f}Y)$
$+\sum\nolimits_{i=1}^{s}\eta^{i}(X)\,g(R_{\xi_{i}\,,{f}Z}V,{f}Y)+(1/2)\,\delta({f}X,Z,V,{f}Y).$ (60)
$+\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)-\eta^{i}(Y)\,\eta^{j}(Z)\,g(h_{i}h_{j}V,QX)$
$+\eta^{i}(Z)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,X)\big]+(1/2)\,\delta({f}X,Z,V,{f}Y).$ (61)
Substituting (3) into (3), we get
$g(R_{X,{f}Z}\,{f}V,Y)-g(R_{X,{f}Z}\,{f}^{2}V,{f}Y)=g(R_{X,Z}\,QV,Y)-g(R_{X,Z}{f}V,{f}Y)$
$-g(R_{\widetilde{Q}X,{f}Z}V,{f}Y)+g(R_{X,{f}Z}\widetilde{Q}V,{f}Y)+g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)+(1/2)\,\delta({f}X,Z,V,{f}Y).$ (62)
Using (45) and (46)–(5), we have
$g(R_{X,Y}{f}V,{f}Z)-g(R_{{f}V,{f}Z}{f}X,{f}Y)=g(R_{V,Z}{f}X,{f}Y)-g(R_{V,Z}QX,QY)$
$+\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}QY,Z)-\eta^{i}(X)\eta^{j}(Z)\,g(h_{i}h_{j}QY,V)$
$+\,\eta^{i}(Y)\,\eta^{j}(Z)\,g(h_{i}h_{j}QX,V)-\eta^{i}(Y)\,\eta^{j}(V)\,g(h_{i}h_{j}QX,Z)\big]$
$-\,(1/2)\,\delta({f}X,{f}Y,Z,V)-(1/2)\,\delta(X,Y,Z,V).$ (63)
Substituting (3), (3), (3) and (3) into (3), and using Lemma 2, we obtain
$g(Q(\nabla_{X}{f})Z,(\nabla_{V}{f})Y)-g(Q(\nabla_{X}{f})Y,(\nabla_{V}{f})Z)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,{f}h_{i}Z)$
$-\eta^{i}(V)\,g((\nabla_{X}{f})Z,{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}{f})Y,{f}h_{i}X)-\eta^{i}(Z)g((\nabla_{X}{f})Y,{f}h_{i}V)$
$+g(h_{i}X,QZ)g(h_{i}V,QY)-g(h_{i}X,Y)\,g(h_{i}V,QZ)\big]+\sum\nolimits_{i,j=1}^{s}\big[2\,\eta^{i}(Z)\eta^{j}(V)g(h_{i}h_{j}Y,X)$
$-\eta^{i}(X)\,\eta^{j}(Z)\,g(h_{i}h_{j}Y,QV)-\eta^{i}(Y)\,\eta^{j}(V)\,g(h_{i}h_{j}X,QZ)\big]+g(R_{X,Z}\,QV,Y)$
$-g(R_{X,Z}{f}V,{f}Y)+g(R_{V,Z}{f}X,{f}Y)-g(R_{V,Z}\,QX,QY)$
$=g(R_{\widetilde{Q}X,{f}Z}V,{f}Y){-}g(R_{X,{f}Z}\widetilde{Q}V,{f}Y){-}g(R_{X,{f}Z}{f}V,\widetilde{Q}Y){-}g(R_{X,Y}{f}V,{f}\widetilde{Q}Z)-g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)$
$-\sum\nolimits_{i=1}^{s}g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)-\sum\nolimits_{i,j=1}^{s}\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)-(1/2)\,\delta({f}X,Z,V,{f}Y)$
$+(1/2)\,\delta({f}X,{f}Y,Z,V)+(1/2)\,\delta(X,Y,Z,V)-(1/4)\,\delta(X,{f}Z,{f}V,Y)-(1/4)\,\delta({f}V,{f}Z,X,Y).$ (64)
Adding (3) to (3), we obtain
$+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,{f}h_{i}Z)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}{f})Y,{f}h_{i}X)$
$-\eta^{i}(Z)g((\nabla_{X}{f})Y,{f}h_{i}V)+g(h_{i}X,QZ)g(h_{i}V,QY)+g(h_{i}V,Y)\,g(h_{i}X,Z)\big]$
$+\sum\nolimits_{i,j=1}^{s}\big[2\,\eta^{i}(Z)\eta^{j}(V)g(h_{i}h_{j}Y,X)-\eta^{i}(X)\eta^{j}(Z)g(h_{i}h_{j}Y,QV)-\eta^{i}(Y)\eta^{j}(V)g(h_{i}h_{j}X,QZ)\big]$
$+2\,g(R_{X,Z}\,V,Y)-2\,g(R_{X,Z}{f}V,{f}Y)-g(R_{V,Z}\,QX,QY)+g(R_{X,Y}V,QZ)$
$-g(R_{X,Y}{f}V,{f}\widetilde{Q}Z)-g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)-\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)$
$+\sum\nolimits_{i=1}^{s}\big[g(h_{i}X,Y)\,g(h_{i}V,\widetilde{Q}Z)-g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)-\nabla_{V}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)\big]$
$-(1/2)\,\delta({f}X,Z,V,{f}Y)+(1/2)\,\delta({f}X,{f}Y,Z,V)+(1/2)\,\delta(X,Y,Z,V)-(1/4)\delta(X,Z,V,Y)$
$-(1/4)\delta(V,Z,X,Y)-(1/4)\,\delta(X,{f}Z,{f}V,Y)-(1/4)\,\delta({f}V,{f}Z,X,Y).$ (65)
$\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,{f}h_{i}Z)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}\,{f})Y,{f}h_{i}X)$
$-\eta^{i}(Y)\,g((\nabla_{Z}\,{f})V,{f}h_{i}X)-\eta^{i}(Z)g((\nabla_{X}{f})Y,{f}h_{i}V)+\eta^{i}(Y)g((\nabla_{Z}{f})X,{f}h_{i}V)$
$-\eta^{i}(X)g((\nabla_{Y}\,{f})V,{f}h_{i}Z)+\eta^{i}(X)g((\nabla_{Z}{f})V,{f}h_{i}Y)\big]$
$+\sum\nolimits_{i,j=1}^{s}\big[2\,\eta^{i}(Z)\,\eta^{j}(V)g(h_{i}h_{j}Y,X)-2\,\eta^{i}(X)\,\eta^{j}(Y)g(h_{i}h_{j}V,Z)]$
$+g(R_{Z,V}\,QX,QY)+g(R_{Z,V}\,QX,Y)-g(R_{X,Y}\,QZ,V)-g(R_{X,Y}\,QZ,QV)$
$=g(R_{\widetilde{Q}X,{f}Z}V,{f}Y)-g(R_{X,{f}Z}{f}V,\widetilde{Q}Y)+g(R_{Z,V}{f}Y,{f}\widetilde{Q}X)$
$-g(R_{X,Y}{f}V,{f}\widetilde{Q}Z)-g(R_{\widetilde{Q}Z,{f}X}Y,{f}V)+g(R_{Z,{f}X}{f}Y,\widetilde{Q}V)$
$+\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(Z)\,\eta^{j}(Y)\,g(h_{i}h_{j}V,\widetilde{Q}X)-\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)\big]$
$+\sum\nolimits_{i=1}^{s}\big[g(h_{i}X,Y)\,g(h_{i}V,\widetilde{Q}Z)-g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)-g(h_{i}Z,V)\,g(h_{i}Y,\widetilde{Q}X)$
$+g({f}h_{i}Z,V)g({f}h_{i}X,\widetilde{Q}Y)\big]+(1/4)\,\delta({f}X,Z,{f}Y,V)+(1/4)\,\delta({f}X,{f}Y,Z,V)$
$+(1/2)\,\delta(X,Y,Z,V)+(1/4)\,\delta(X,{f}Z,{f}V,Y)-(1/4)\,\delta({f}Z,{f}V,X,Y)$
$+\sum\nolimits_{i,j=1}^{s}\big[\nabla_{Y}\big(\eta^{i}(X)\,g(h_{i}Z,\widetilde{Q}V)\big)-\nabla_{V}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)\big].$ (66)
Putting $\xi_{k}$ in place of $V$ in (3), then using Lemma 2, (12), $\nabla_{\xi_{k}}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)=0$, Remark 1 and (8), we get
$\qquad g((\nabla_{X}{f})Y,{f}h_{k}Z)=g((\nabla_{X}{f})Z,{f}h_{k}Y)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Z)g(h_{k}Y,h_{i}X)-\eta^{i}(Y)g(Qh_{k}Z,h_{i}X)$
$-\eta^{i}(Z)g(h_{k}\widetilde{Q}Y,h_{i}X){-}\eta^{i}(X)g(h_{i}h_{k}Y,\widetilde{Q}Z)\big]{+}g(R_{\xi_{k},Z}QX,QY){+}g(R_{\xi_{k},Z}QX,Y){-}2g(R_{\xi_{k},QZ}X,Y).$
Simplifying the above using (1) and (45), we have
$g((\nabla_{Y}{f})X,{f}h_{k}Z)+g((\nabla_{X}{f})Z,{f}h_{k}Y)=-\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Z)g(h_{k}Y,h_{i}X)$
$\quad-\eta^{i}(Y)g(Qh_{k}Z,h_{i}X)-\eta^{i}(Z)g(h_{k}\widetilde{Q}Y,h_{i}X)+2\eta^{i}(X)g(h_{i}Y,\widetilde{Q}h_{k}Z)\big].$ (67)
Swapping $X$ and $Y$ in the above equation and adding it back to (3), we have
$g((\nabla_{Y}f)Z,fh_{k}X)+g((\nabla_{X}f)Z,fh_{k}Y)=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(X)g(h_{i}Y,Qh_{k}Z)+\eta^{i}(Y)g(h_{i}X,Qh_{k}Z)$
$-2\eta^{i}(Z)g(h_{i}X,h_{k}Y)+2\eta^{i}(Z)g(h_{i}X,\widetilde{Q}h_{k}Y)-2\eta^{i}(X)g(h_{i}Y,\widetilde{Q}h_{k}Z)-2\eta^{i}(Y)g(h_{i}X,\widetilde{Q}h_{k}Z)\big].$ (68)
By swapping $Z$ and $X$ in (3), and adding the resulting equation and (3), we get
$2\,g((\nabla_{Y}f)X,fh_{k}Z)=2\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)g(h_{k}X,h_{i}Z)-\eta^{i}(X)g(h_{i}Z,h_{k}Y)\big].$
Finally, simplifying the above using (8) and the definition of $\widetilde{Q}$, we get the desired expression (49). ∎
4 Main results
In this section, we prove a splitting result (Theorem 1) for weak nearly $\cal C$}-manifolds satisfying some additional conditions such as (9) and (13). In Theorem 2, we characterize $(4+s)$-dimensional weak nearly $\cal C$}-manifolds. The consequences of Theorems 1 and 2 (without the requirement of (9) and (13)) present new results for nearly $\cal C$}-manifolds.
The following result generalizes [18, Proposition 4.1] (and [14, Proposition 4.2]) to the case $s>1$.
Proposition 1.
For a weak nearly $\cal C$}-manifold with conditions (5), (6), (9) and (13), the eigenvalues (and their multiplicities) of the symmetric operators $h_{i}h_{j}$ and $h_{i}^{2}$ are constant for all $1\leq i\neq j\leq s$. Moreover, each $h_{i}h_{j}$ and $h_{i}^{2}$ shares the same eigen-frame.
Proof.
From (45) and Lemma 5 we obtain
$(\nabla_{X}\,h_{i}h_{j})Y=h_{i}(\nabla_{X}\,h_{j})Y+(\nabla_{X}\,h_{i})h_{j}Y=\sum\nolimits_{k=1}^{s}\big[g(X,h_{k}h_{i}h_{j}Y)\,\xi_{k}-\eta^{j}(Y)\,h_{i}h_{j}h_{k}X\big].$ (69)
@@ 4 Main results
which implies that $X(\mu)=0$ for all $X\in\mathfrak{X}_{M}$. When $j=i$, we get a similar claim for $h_{i}^{2}$.
Next, by Lemma 6, the operators $h_{i}$ and $h_{j}$ commute for all $i,j$, which implies that $h_{i}^{2}$ and $h_{j}^{2}$ also commute. Each $h_{i}^{2}$ is self-adjoint, and therefore, diagonalizable.
It is well known that two diagonalizable operators on a finite-dimensional Euclidean space are simultaneously diagonalizable if and only if they commute. This fact extends by mathematical induction to any finite family of commuting diagonalizable operators. By the above, there exists an orthonormal basis $\{e_{k}\}$ of the tangent space $TM$ such that $h_{i}^{2}e_{k}=\lambda_{i,k}\,e_{k}$ for all $1\leq i\leq s$ and $k$. ∎
@@
where $\lambda_{i,j}$ is a positive real number and $\lambda_{i,j}\neq\lambda_{i,k}$ for $j\neq k$. If $X\neq 0$ is an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$, then $X,{f}X,h_{i}X$ and $h_{i}\,{f}X$ are orthogonal nonzero eigenvectors of $h_{i}^{2}$ with the same eigenvalue $\lambda^{2}_{i,j}$. Since $h_{i}(\xi_{j})=0$ for all $1\leq i,j\leq s$, the Reeb distribution $\ker f=\operatorname{span}\{\xi_{1},\dots,\xi_{s}\}$ is contained in the kernel of $h_{i}$. Consequently, the eigenvalue 0 of $h_{i}^{2}$ has multiplicity $2p_{i}+s$ for some integer $p_{i}\geq 0$, associated with each operator $h_{i}^{2}$. Denote by
@@
$d\eta^{i}(X,Y)$ $=g(\nabla_{X}Y,\xi_{i})+g(Y,\nabla_{X}\xi_{i})-g(\nabla_{Y}X,\xi_{i})-g(X,\nabla_{Y}\xi_{i})-g([X,Y],\xi_{i})$
$=g(Y,h_{i}X)-g(X,h_{i}Y)=2\,g(h_{i}X,Y)\quad(1\leq i\leq s).$ (71)
@@
The following result generalizes [18, Proposition 4.3] (and [14, Proposition 4.1]) to the case $s>1$.
@@
By (2) and the condition ${\bf h}\equiv 0$, we have $d\eta^{i}=0$ for all $i$. We have $0=d\eta^{i}(X,Y)=-\eta^{i}([X,Y])$ for all $i$ and $X,Y\in{\cal D}$, which indicates that $X,Y]\in{\cal D}$; thus the distribution $\cal D$} is integrable. Integral manifolds of $\cal D$} are totally geodesic as $g(\nabla_{X}\,Y,\xi_{i})=-g(Y,h_{i}X)=0$ for $X,Y\in{\cal D}$ and $1\leq i\leq s$. Since $\operatorname{ker}f$ is integrable, totally geodesic and flat (see Lemma 1), by de Rham Decomposition Theorem, e.g., [12], the manifold is locally the Riemannian product $\mathbb{R}^{s}\times\bar{M}$. The weak nearly $\cal C$}-structure on $M$ induces on $\bar{M}$ a weak nearly Kähler structure.
@@
The following result generalizes [18, Proposition 4.2] (and [14, Proposition 4.3]) to the case $s>1$.
@@
    ${\cal D}_{i,0}\neq 0$ and distributions $\operatorname{ker}f\oplus{\cal D}_{i,0}$ are integrable with totally geodesic leaves;
@@
    if the exterior derivatives of the coframe $\{\eta^{i}\}$ have a common kernel, then distributions ${\cal D}_{i,0}$ coincide, are integrable with totally geodesic leaves, and the induced structure on each leaf of ${\cal D}_{i,0}$ is a weak nearly Kähler structure $(\bar{f},\bar{g})$ with the property $\bar{\nabla}(\bar{f}^{\,2})=0$;
@@
$h_{i}^{2}(\nabla_{X}\,Y)=-\lambda^{2}_{i,j}\,\nabla_{X}Y-(\nabla_{X}\,h_{i}^{2})Y=-\lambda^{2}_{i,j}\,\nabla_{X}Y+\lambda^{2}_{i,j}\sum\nolimits_{k=1}^{s}g(X,h_{i}Y)\,\xi_{k}.$
@@
$\nabla_{X}Y=-\widetilde{Q}\,\nabla_{X}Y-{f}^{2}\nabla_{X}Y+\sum\nolimits_{k=1}^{s}\eta^{k}(\nabla_{X}Y)\,\xi_{k},$
@@
Since $\mathcal{D}_{i,0}\neq 0$, we get $h_{i}^{\,2}(\nabla_{X}Y)=0$ for all $X,Y\in\mathcal{D}_{i,0}$. In addition, we compute $g(\nabla_{X}Y,\xi_{k})=-g\big(Y,\nabla_{X}\xi_{k}\big)=-g(Y,h_{k}X)$, which vanishes if $\operatorname{ker}h_{i}=\operatorname{ker}h_{k}$ for all $1\leq i,k\leq s$. By (2), this condition is equivalent to the fact that the exterior derivatives of the coframe $\{\eta^{i}\}$ have a common kernel. Thus, each distribution $D_{i,0}$ defines a totally geodesic foliation if and only if the exterior derivatives of the coframe $\{\eta^{i}\}$ have a common kernel. By (8) and (18), the leaves of $D_{i,0}$ are f{f}-invariant and $Q$-invariant. Thus, the weak nearly $\cal C$}-structure on $M$ with conditions (5), (6), (9) and (13) induces a weak nearly Kähler structure $(\bar{f},\bar{g})$ on each leaf of $D_{i,0}$ with the property $\bar{\nabla}(\bar{f}^{\,2})=0$, where $\bar{\nabla}$ is the Levi-Civita connection of $\bar{g}$. This proves $(iii)$.
To establish $(iv)$, it suffices to verify that for $X\in{\cal D}_{i,j}$ and $Y\in{\cal D}_{i,k}$ with $j\neq k$, one has $\nabla_{X}Y\in\ker f\oplus{\cal D}_{i,1}\oplus\cdots\oplus{\cal D}_{i,r}$. Similarly to part $(i)$, we obtain $h_{i}^{2}(\nabla_{X}Y)=-\lambda_{i,k}\,\nabla_{X}Y+\lambda_{i,k}\sum_{j=1}^{s}g(X,h_{j}Y)\,\xi_{j}$; hence, $h_{i}^{2}(f^{2}\nabla_{X}Y)=-\lambda_{i,k}\,f^{2}\nabla_{X}Y$. Thus $f^{2}\nabla_{X}Y\in{\cal D}_{i,k}$, and we calculate $\widetilde{Q}\nabla_{X}Y\in{\cal D}_{i,k}$. Combining these observations, we may write

### v2

@@ 3 Main results
The first part follows from the commutativity and skew-symmetry of the operators $h_{i}\ (1\leq i\leq s)$.
From the above and Lemma 5 we obtain
$(\nabla_{X}\,h_{i}h_{j})Y=h_{i}(\nabla_{X}\,h_{j})Y+(\nabla_{X}\,h_{i})h_{j}Y=\sum\nolimits_{k}\big[g(X,h_{k}h_{i}h_{j}Y)\,\xi_{k}-\eta^{j}(Y)\,h_{i}h_{j}h_{k}X\big].$ (15)
@@
which implies that $X(\mu)=0$ for all $X\in\mathfrak{X}_{M}$. When $j=i$, we get a similar result for $h_{i}^{2}$.
Next, by the above, the operators $h_{i}$ and $h_{j}$ commute for all $i,j$, which implies that $h_{i}^{2}$ and $h_{j}^{2}$ also commute. Each $h_{i}^{2}$ is self-adjoint, and therefore, diagonalizable.
It is well known that a finite family of diagonalizable operators on a finite-dimensional Euclidean space is simultaneously diagonalizable if and only if the opeators commute. By the above, there exists an orthonormal basis $\{e_{k}\}$ of $TM$ such that $h_{i}^{2}e_{k}=\lambda_{i,k}\,e_{k}$ for all $1\leq i\leq s$ and $k$. ∎
@@
where $r_{i}\geq 1$, $\lambda_{i,j}$ is a positive real number and $\lambda_{i,j}\neq\lambda_{i,k}$ for $j\neq k$. If $X\neq 0$ is an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$, then $X,{f}X,h_{i}X$ and $h_{i}\,{f}X$ are orthogonal nonzero eigenvectors of $h_{i}^{2}$ with the same eigenvalue $\lambda^{2}_{i,j}$. Since $h_{i}(\xi_{j})=0$ for all $1\leq i,j\leq s$, the Reeb distribution $\ker f=\operatorname{span}\{\xi_{1},\dots,\xi_{s}\}$ is contained in the kernel of $h_{i}$. Consequently, the eigenvalue 0 of $h_{i}^{2}$ has multiplicity $2p_{i}+s$ for some integer $p_{i}\geq 0$, associated with each operator $h_{i}^{2}$. Denote by
@@
$d\eta^{i}(X,Y)=2\,g(h_{i}X,Y)\quad(1\leq i\leq s).$ (17)
@@
The following result generalizes [20, Proposition 4.3] (and [16, Proposition 4.1]) to the case $s>1$.
@@
By (17) and the condition ${\bf h}\equiv 0$, we have $d\eta^{i}=0$ for all $i$. We have
$0=d\eta^{i}(X,Y)=-\eta^{i}([X,Y])\quad(1\leq i\leq s,\ X,Y\in{\cal D}),$
which indicates that $X,Y]\in{\cal D}$; thus the distribution $\cal D$} is integrable. Integral manifolds of $\cal D$} are totally geodesic as
$g(\nabla_{X}\,Y,\xi_{i})=-g(Y,h_{i}X)=0\quad(1\leq i\leq s,\ X,Y\in{\cal D}).$
Since $\operatorname{ker}f$ is integrable, totally geodesic and flat (see Lemma 1), by de Rham Decomposition Theorem, the manifold is locally the Riemannian product $\mathbb{R}^{s}\times\bar{M}$. The weak nearly $\cal C$}-structure on $M$ induces on $\bar{M}$ a weak nearly Kähler structure.
@@
The following result generalizes [20, Proposition 4.2] (and [16, Proposition 4.3]) to the case $s>1$.
@@
    ${\cal D}_{i,0}\neq 0$ for all $i$, and distributions $\operatorname{ker}f\oplus{\cal D}_{i,0}$ are integrable with totally geodesic leaves;
@@
    if the exterior derivatives of the coframe $\vec{\eta}$ have a common kernel, then distributions ${\cal D}_{i,0}$ coincide, are integrable with totally geodesic leaves, and the induced structure on each leaf of ${\cal D}_{i,0}$ is a weak nearly Kähler structure $(\bar{f},\bar{g})$ with the property $\bar{\nabla}(\bar{f}^{\,2})=0$;
@@
$h_{i}^{2}(\nabla_{X}\,Y)=-\lambda^{2}_{i,j}\,\nabla_{X}Y-(\nabla_{X}\,h_{i}^{2})Y=-\lambda^{2}_{i,j}\,\nabla_{X}Y+\lambda^{2}_{i,j}\sum\nolimits_{k}g(X,h_{i}Y)\,\xi_{k}.$
@@
$\nabla_{X}Y=-\widetilde{Q}\,\nabla_{X}Y-{f}^{2}\nabla_{X}Y+\sum\nolimits_{k}\eta^{k}(\nabla_{X}Y)\,\xi_{k},$
@@
Since $\mathcal{D}_{i,0}\neq 0$, we get $h_{i}^{\,2}(\nabla_{X}Y)=0$ for all $X,Y\in\mathcal{D}_{i,0}$. In addition, we compute
$g(\nabla_{X}Y,\xi_{k})=-g\big(Y,\nabla_{X}\xi_{k}\big)=-g(Y,h_{k}X),$
which vanishes if $\operatorname{ker}h_{i}=\operatorname{ker}h_{k}$ for all $1\leq i,k\leq s$. By (17), this condition is equivalent to the fact that the exterior derivatives of the coframe $\vec{\eta}$ have a common kernel. Thus, each distribution $D_{i,0}$ defines a totally geodesic foliation if and only if the exterior derivatives of the coframe $\vec{\eta}$ have a common kernel. By (9) and Lemma 2, the leaves of $D_{i,0}$ are f{f}-invariant and $Q$-invariant. Thus, the weak nearly $\cal C$}-structure on $M$ with conditions (6), (7), (10) and (13) induces a weak nearly Kähler structure $(\bar{f},\bar{g})$ on each leaf of $D_{i,0}$ with the property $\bar{\nabla}(\bar{f}^{\,2})=0$, where $\bar{\nabla}$ is the Levi-Civita connection of $\bar{g}$. This proves $(iii)$.
To establish $(iv)$, it suffices to verify that for $X\in{\cal D}_{i,j}$ and $Y\in{\cal D}_{i,k}$ with $j\neq k$, one has $\nabla_{X}Y\in\ker f\oplus{\cal D}_{i,1}\oplus\cdots\oplus{\cal D}_{i,r}$. Similarly to part $(i)$, we obtain
$h_{i}^{2}(\nabla_{X}Y)=-\lambda_{i,k}\,\nabla_{X}Y+\lambda_{i,k}\sum\nolimits_{j}g(X,h_{j}Y)\,\xi_{j};$
hence, $h_{i}^{2}(f^{2}\nabla_{X}Y)=-\lambda_{i,k}\,f^{2}\nabla_{X}Y$. Thus $f^{2}\nabla_{X}Y\in{\cal D}_{i,k}$, and we calculate $\widetilde{Q}\nabla_{X}Y\in{\cal D}_{i,k}$. Combining these observations, we may write